# Why Does Little Robustness Help? A Further Step Towards Understanding Adversarial Transferability

### v7

@@ 3.1 Data Augmentation Mechanisms
Mixup (MU) [84] trains a model on the convex combination of randomly selected sample pairs, i.e., the mixed image $\tilde{x}_{i,j}=bx_{i}+(1-b)x_{j}$ and the corresponding label $\tilde{y}_{i,j}=by_{i}+(1-b)y_{j}$, where $(x_{i},y_{i}),(x_{j},y_{j})\in\mathbf{S}$. Here $b\in[0,1]$ is a random variable drawn from the Beta distribution $\operatorname{Beta}(1,1)$. Additionally, we use a probability parameter $\mathbf{p}$ to control how much $\mathcal{D}^{\prime}$ shifts away from $\mathcal{D}$. Specifically, as $\mathbf{p}\rightarrow 0$, the augmented dataset will be identical to $\mathbf{S}$; as $\mathbf{p}\rightarrow 1$, all the samples will be interpolated. Thus, at a high level, $\mathbf{p}$ controls how much $\mathcal{D}^{\prime}$ shifts away from $\mathcal{D}$. We train augmented models for MU with $\mathbf{p}\in[0.1,0.3,0.5,0.7,0.9]$.
@@
Cutmix (CM[82] also augments both images and labels by mixing samples. The difference is patches are cut and pasted among training samples. A mixed sample $\tilde{x}_{i,j}=\mathbf{M_{b}}\odot x_{i}+(\mathbf{1}-\mathbf{M_{b}})\odot x_{j}$ and its label $\tilde{y}_{i,j}$ are also obtained from two samples $(x_{i},y_{i}),(x_{j},y_{j})\in\mathbf{S}$, where $\mathbf{M_{b}}$ is a binary matrix with the same size of $x_{i}$ to indicate the location for cutting and pasting, $\mathbf{1}$ is a matrix filled with 1, and $b$ is drawn from $\operatorname{Beta}(1,1)$. $\tilde{y}_{i,j}$ is obtained similarly as MU. Similar to in MU, we use a probability parameter $\mathbf{p}$ to control the augmentation magnitude and train augmented models for CM with $\mathbf{p}\in[0.1,0.3,0.5,0.7,0.9]$.
Cutout (CO) [10] augments the dataset by masking out random regions of images with a size less than $M\times M$. In the real implementation, the mask-out region and its size for each image are not deterministic. A bigger $M$ value indicates it is more likely to mask out a bigger region. Hence, we utilize the value of $M$ only to regulate the extent of augmentation since it probabilistically governs the augmentation magnitude, similar to the role of $\mathbf{p}$ in MU and CM. We select $M\in[8,12,16,20,24]$ and $M\in[80,100,120,140,160]$ for CIFAR-10 and ImageNette, respectively.
Label smoothing (LS) [62] augments the dataset by replacing the hard labels $\mathcal{L}$ with soft continuous labels in probability distribution. Specifically, the probability of the ground-truth class is $1-p$, and the probability of other classes will be uniformly assigned to $p/(m-1)$, where $p\in[0,1)$ and $m$ denotes the number of classes. We train models for LS with $p\in[0.1,0.2,0.3,0,4,0.5]$.
@@ 3.3 Trade-off Under Data Augmentation
First, LS exhibits a monotonous benefit on smoothness on both datasets due to its implicit effect on reducing the gradient norm during training [62]. However, the tables show that the ASRs of LS are always lower than those of ST, implying this benefit may not completely offset the negative effect of largely degraded similarity in LS. Additionally, there is no monotonicity in ASRs in LS, aligning with the always combating negative and positive effects. In contrast, other augmentations do not exhibit a consistent tendency for smoothness on both datasets. Specifically, in CIFAR-10, CM, MU, and CO mostly degrade the smoothness while the degradation is not strictly monotonous. The situation is even more chaotic in ImageNette, where the variances of smoothness are relatively large. However, the ASRs do reveal some patterns. For CM and MU, their ASRs degrade from $\tau=1$ to $\tau=5$, and the ASRs of CO are generally better than MU and CM. Notably, CO performs slightly better than ST under a few cases, especially when $\tau=1$. This agrees with the fact that CO has slightly better smoothness and less severely degraded similarity. Conclusively, the trade-off under data augmentations is quite complex, and no single augmentation can produce good surrogates.
@@ 4.1.1 Regularization in the Input Space
Based on the implication that regularizing input space smoothness benefits transferability shown in Sec. 2.2, one should directly optimize the loss surface curvature. However, it is challenging to realize it since the computationally expensive second-order derivative is involved, i.e., $\nabla_{x}^{2}\ell(f_{\theta}(x))$. Fortunately, deep learning and optimization theory [16] allow us to use the first-order derivative to approximate it. We thus choose two well-known first-order gradient regularizations as follows:
@@
where $\lambda_{ir}$ controls the regularization magnitude. IR has been demonstrated improving the generalization and interpretability of DNNs [16, 51]. Here we show that IR can also improve transferability. Regularizing $\|\nabla_{x}\ell(f(x))\|_{2}$ leads to a smaller curvature $\sigma(\nabla_{x}^{2}\ell(x))$ because the spectral norm of a matrix is upper-bounded by the Frobenius norm, i.e., $\sigma(\nabla_{x}^{2}\ell(x))\leq\|\nabla_{x}^{2}\ell(x)\|_{F}\approx\|\nabla_{x}\ell(f(x))^{T}\nabla_{x}\ell(f(x))\|_{F}\leq\|\nabla_{x}\ell(f(x))\|^{2}$. Here we approximate the second-order derivative using the first-order derivative as suggested in [43, 28]. Consequently, smaller $\sigma(\nabla_{x}^{2}\ell(x))$ indicates better smoothness, thereby improving transferability.
@@
JR has been also correlated with the generalization and robustness [28, 80, 4, 25, 64]. Since JR has a similar form with IR, it is reasonable to presume that they have a similar effect. Formally, we prove that when $\|\nabla_{x}f_{\theta}(x_{i})\|_{F}\rightarrow 0$, we have $\|\nabla_{x}\ell(f_{\theta}(x_{i}))\|\rightarrow 0$ using cross-entropy loss as an example (see Proposition 1). This implies that JR also has a positive effect on model smoothness.
@@ 4.1.2 Regularization in the Weight Space
Recently, researchers paid attention to the gradient regularization on $\theta$, i.e., the parameters (weights) of the network. A recent study [11] proved that the gradient regularizing pressure on the weight space $\|\nabla_{\theta}f_{\theta}(x)\|_{F}$ can transfer to the input space $\|\nabla_{x}f_{\theta}(x)\|_{F}$. Consequently, regularizing gradients w.r.t. the weight space could also implicitly promote the smoothness in the input space. We next explore two representative weight gradient regularization methods as follows:
@@
where $L$ is the original objective function. Similar to IR and JR, we can directly apply the double-propagation technique to implement the regularization term, using the automatic differentiation framework provided by modern ML libraries like PyTorch [47]. ER has been empirically studied on improving generalization [3, 55, 20], while we investigate it from the transferability perspective.
Sharpness-aware minimization (SAM): SAM also seeks flat solution in the weight space. Specifically, it intends to minimize the original training loss and the worst-case sharpness $L_{sam}(\theta)=L(\theta)+[\max_{\|\boldsymbol{\hat{\theta}-\theta}\|_{2}\leq\rho}L(\hat{\theta})-L(\theta)]=\max_{\|\boldsymbol{\hat{\theta}-\theta}\|_{2}\leq\rho}L(\hat{\theta})$, where $\rho$ denotes the radius to search for the worst neighbor $\hat{\theta}$. As the exact worst neighbor is difficult to track, SAM uses the gradient of ascent direction neighbor for the update after twice approximations:
$\nabla_{\theta}L_{sam}(\theta)\approx\nabla_{\theta}L(\theta+\rho\frac{\nabla_{\theta}L(\theta)}{||\nabla_{\theta}L(\theta)||}).$
Recent researches establish SAM as a special kind of gradient normalization [85, 29], where $\rho$ represents the regularization magnitude. SAM has been extensively studied recently for its remarkable performance on generalization [18, 2, 6, 90, 70, 17, 1]. In this paper, for the first time, we substantiate that SAM also features outstanding superiority in boosting transferability.
@@ 4.2 Gradient Regularization Promotes Smoothness
We train multiple regularized models for each regularization and use coarse-grained parameter intervals [51, 18, 55] and restrict the largest regularized magnitude to avoid excessive regularization such that the accuracy of these models are above 90%. This allows for a fair comparison of each regularization method in an acceptable accuracy range. We choose 5 parameters in each interval through random grad and binary choices (See Tab. XI).
@@
Figs. 9 and 10 depict the model smoothness. We can observe gradient regularizations generally result in smoother models (i.e., smaller values) than the baseline where the regularization magnitude is 0 (ST), and large magnitudes correspond to better smoothness. Moreover, input gradient regularizations yield much better smoothness than weight gradient regularizations, particularly on CIFAR-10. This indicates that the transfer effect from weight space to input space suggested in [11] does exist, yet is somewhat marginal and cannot outperform direct regularizations in input space.
@@ 4.3 Trade-off Under Gradient Regularization
One might presume that input gradient regularization (JR, IR) should clearly outperform weight gradient regularization (SAM, ER) in terms of transferability, which agrees with the conclusion that less complexity should exhibit better transferability in Demontis et al. [9], given that smoothness is positively correlated to the model complexity (as analyzed in Sec. 6, Q3). However, our experimental results counter this:
@@
Therefore, these inconsistent results do not allow us to conclude that better smoothness always leads to better or worse transferability. This finding not only weakens the conclusion in Demontis et al. [9], but also suggests an overall examination of these training mechanisms for superior surrogates, as they may weigh differently on similarity.
@@ 5.2.1 Incorporating AE generation strategies
Exploiting the generation process of AEs is a popular approach to improve transferability. The most representative methods include momentum iterative (MI) [12] and diversity iterative (DI) [76], and DI is frequently combined with MI and serves as a stronger baseline—diversity iterative with momentum (DIM). Generally, these designs are interpreted as constructing adversarial examples that are less likely to fall into a poor local optimum of the given surrogate. Thus they adjust the update direction with momentum to escape it or apply multiple transformations on inputs to mimic an averagely flat landscape. We conduct experiments to verify that obtaining a better surrogate, which features less and wider local optima (analyzed in Sec. 6, Q3) and general adversarial direction, can further enhance these designs by promoting a larger lower bound.
@@ 5.2.2 Incorporating ensemble strategies
Using multiple surrogates is widely believed to update adversarial examples in a more general direction that benefits transferability. Recently, [23] proposed constructing the ensemble from a large geometric vicinity (LGV), i.e., first fine-tuning a normally trained model with a high constant learning rate to obtain a set of intermediate models that belongs to a large vicinity, then optimizing AEs iteratively against all these models. We conduct experiments to verify whether SAM and SAM&JR can further enhance this ensemble and analyze how their intrinsic mechanisms correlate with LGV. Specifically, we fine-tune models well-trained with SGD [^3] We refer SGD here to training models without any augmentation or regularization mechanisms, not merely the particular optimizer itself., SAM, and SAM&JR on CIFAR-10 using SGD, SAM, and SAM&JR with a constant learning rate of 0.05.
In Tab. V, the transfer results show that LGV with SAM, SAM&JR can always yield superior ASRs than standard SGD, indicating that the underlying property of SAM and JR allows them to find solutions that benefit transferability. More intriguingly, fine-tuning models obtained by SAM and SAM&JR with SGD always brings non-negligible negative impacts on transferability. These suggest solutions well-trained with SAM and SAM&JR lie in the basin of attractions featuring better transferability, and a relatively high constant learning rate with SGD will arbitrarily minimize the training loss while leaving the basin of attractions. Note that a relatively high constant learning rate is necessary for a large vicinity [23], thus we can infer that a large vicinity may not be a determining factor for transferability since a single surrogate can outperform an ensemble of poor surrogates (see SAM&JR and ST+$\rm{LGV_{\text{SGD}}}$). Instead, a dense subspace whose geometry is intrinsically connected to transferability is more desirable.
@@ 5.2.3 Incorporating other strategies
We take AE generation and ensemble strategies, which are the most well-explored methods, to exemplify the universality of better surrogates. We are aware of other factors that could also contribute to transferability, such as loss object modification [86], architecture selection [72], $l_{2}$-norm consideration [42], unrestricted generation [26, 87], patch-based AEs [27]. We leave it to our future work to investigate whether they can be mildly integrated with our design.
@@ 6 Analyses and Discussions
Table VI: The Pearson correlation coefficients ($r$) between ASR and accuracy ($Acc$), fooling probability ($FP$), model smoothness ($MS$), and gradient similarity ($GS$). For the coefficient of determination ($R^{2}$) in OLS, $R_{1}^{2}$ depicts the joint effect of $MS$ and $GS$ on ASR and $R_{2}^{2}$ depicts the joint effect of these 4 factors.
@@
Q2: What if the target model is not trained on $\mathcal{D}$? So far, we presume the target models are trained within the normal data distribution $\mathcal{D}$ to lay out our analysis and conclusions, which is also set by default in Demontis et al. [9]. To verify the generality of the obtained conclusion, we deliver the results here when the data distribution restrictions of target model $\mathcal{G}$ is relaxed. First, we generalize Hypothesis 1 as:
@@
Q3: What is the relationship between model smoothness and model complexity? Model complexity has been investigated in Demontis et al. [9] to figure out whether it correlates with transferability, which is defined as the variability of the loss landscape, deriving from the number of local optima of the surrogates. Here we correlate the number of local optima with model smoothness. It is obvious that a high-complexity model featuring multiple local optima will cause a larger variance to the loss landscape of AE optimization. Formally, for a feature $x$, the fewer optima in its $\epsilon$-neighbor, the smaller the variance would be. Thus, supposing the loss function used to optimize adversarial points is continuously differentiable, by promoting the smoothness, we can expect fewer optima in the finite $\epsilon$-neighbor space since all the local optima are wider. Consequently, model smoothness negatively correlates with model complexity, i.e., a smoother model indicates lower complexity.
@@
Q4: Does a smaller model generalization gap imply higher adversarial transferability? It is natural to raise this question as the idea of adversarial transferability in a way resembles the concept of model generalization. The former captures the effectiveness of AEs against unseen models, while the latter evaluates the model’s ability to adapt properly to previously unseen data. A concurrent work [68] attempts to show the positive relationship between model generalization and adversarial transferability, by applying a spectral normalization method constraining the $L_{2}$-operator norm of the surrogate’s weight matrix to empirically demonstrate the increase of both transferability and generalization.
Nevertheless, we emphasize that we cannot over-generalize the alignment between generalization and transferability. It is easy to find some cases that boost generalization yet degrades transferability (e.g., through damaging similarity). For example, the results in Tabs. II, III, XIII and X have shown augmentations that boost generalization could sacrifice transferability. More typically, the strong data augmentations that yield the smallest generalization gaps (see Tab. XII) perform the worst in transfer attacks. In fact, the product of spectral norm of model’s layer-wise weight matrices is a loose upper bound of $\|\nabla_{x}f_{\theta}(x)\|_{2}$ [52, 80]. When $\|\nabla_{x}f_{\theta}(x)\|_{2}\rightarrow 0$, we have $\|\nabla_{x}f_{\theta}(x)\|_{F}\rightarrow 0$. Utilizing Proposition 1, we also have $\|\nabla_{x}\ell(f_{\theta}(x),y)\|\rightarrow 0$. This suggests that spectral normalization actually benefits transferability by promoting smoothness.
Q5: What is restraining the adversarial transferability of ViT? Recently, ViT [15] receives notable attention for its outstanding performance on various tasks. Meanwhile, some works [54, 41] have demonstrated that ViT models exhibit lower adversarial transferability than CNNs when used as surrogates, and a recent work [46] argues that this is because the current attacks are not strong enough to fully exploit the transferability potential of ViT. Complementarily, we provide a different viewpoint that the transferability of ViT may have been restrained by its own training paradigm. ViT is known to have poor trainability [7, 6], and thus generally requires large-scale pre-training and fine-tuning with strong data augmentations. This training paradigm is used by default since its appearance. However, as we discussed in this paper, changing the data distribution could have negative impacts on similarity. Using data augmentation is plausible for accuracy performance, yet not ideal for transfer attacks.

### v8

@@ 3.1 Data Augmentation Mechanisms
Mixup (MU) [zhang2018mixup] trains a model on the convex combination of randomly selected sample pairs, i.e., the mixed image $\tilde{x}_{i,j}=bx_{i}+(1-b)x_{j}$ and the corresponding label $\tilde{y}_{i,j}=by_{i}+(1-b)y_{j}$, where $(x_{i},y_{i}),(x_{j},y_{j})\in\mathbf{S}$. Here $b\in[0,1]$ is a random variable drawn from the Beta distribution $\operatorname{Beta}(1,1)$. Additionally, we use a probability parameter $\mathbf{p}$ to control how much $\mathcal{D}^{\prime}$ shifts away from $\mathcal{D}$. Specifically, as $\mathbf{p}\rightarrow 0$, the augmented dataset will be identical to $\mathbf{S}$; as $\mathbf{p}\rightarrow 1$, all the samples will be interpolated. Thus, at a high level, $\mathbf{p}$ controls how much $\mathcal{D}^{\prime}$ shifts away from $\mathcal{D}$. We train augmented models for MU with $\mathbf{p}\in[0.1,0.3,0.5,0.7,0.9]$.
@@
Cutmix (CM) [yun2019cutmix] also augments both images and labels by mixing samples. The difference is patches are cut and pasted among training samples. A mixed sample $\tilde{x}_{i,j}=\mathbf{M_{b}}\odot x_{i}+(\mathbf{1}-\mathbf{M_{b}})\odot x_{j}$ and its label $\tilde{y}_{i,j}$ are also obtained from two samples $(x_{i},y_{i}),(x_{j},y_{j})\in\mathbf{S}$, where $\mathbf{M_{b}}$ is a binary matrix with the same size of $x_{i}$ to indicate the location for cutting and pasting, $\mathbf{1}$ is a matrix filled with 1, and $b$ is drawn from $\operatorname{Beta}(1,1)$. $\tilde{y}_{i,j}$ is obtained similarly as MU. Similar to in MU, we use a probability parameter $\mathbf{p}$ to control the augmentation magnitude and train augmented models for CM with $\mathbf{p}\in[0.1,0.3,0.5,0.7,0.9]$.
Cutout (CO) [devries2017improved] augments the dataset by masking out random regions of images with a size less than $M\times M$. In the real implementation, the mask-out region and its size for each image are not deterministic. A bigger $M$ value indicates it is more likely to mask out a bigger region. Hence, we utilize the value of $M$ only to regulate the extent of augmentation since it probabilistically governs the augmentation magnitude, similar to the role of $\mathbf{p}$ in MU and CM. We select $M\in[8,12,16,20,24]$ and $M\in[80,100,120,140,160]$ for CIFAR-10 and ImageNette, respectively.
Label smoothing (LS) [szegedy2016rethinking] augments the dataset by replacing the hard labels $\mathcal{L}$ with soft continuous labels in probability distribution. Specifically, the probability of the ground-truth class is 1p1-p, and the probability of other classes will be uniformly assigned to p/(m1)p/(m-1), where $p\in[0,1)$ and $m$ denotes the number of classes. We train models for LS with $p\in[0.1,0.2,0.3,0,4,0.5]$.
@@ 3.3 Trade-off Under Data Augmentation
First, LS exhibits a monotonous benefit on smoothness on both datasets due to its implicit effect on reducing the gradient norm during training [szegedy2016rethinking]. However, the tables show that the ASRs of LS are always lower than those of ST, implying this benefit may not completely offset the negative effect of largely degraded similarity in LS. Additionally, there is no monotonicity in ASRs in LS, aligning with the always combating negative and positive effects. In contrast, other augmentations do not exhibit a consistent tendency for smoothness on both datasets. Specifically, in CIFAR-10, CM, MU, and CO mostly degrade the smoothness while the degradation is not strictly monotonous. The situation is even more chaotic in ImageNette, where the variances of smoothness are relatively large. However, the ASRs do reveal some patterns. For CM and MU, their ASRs degrade from $\tau=1$ to $\tau=5$, and the ASRs of CO are generally better than MU and CM. Notably, CO performs slightly better than ST under a few cases, especially when $\tau=1$. This agrees with the fact that CO has slightly better smoothness and less severely degraded similarity. Conclusively, the trade-off under data augmentations is quite complex, and no single augmentation can produce good surrogates.
@@ 4.1.1 Regularization in the Input Space
Based on the implication that regularizing input space smoothness benefits transferability shown in Sec. 2.2, one should directly optimize the loss surface curvature. However, it is challenging to realize it since the computationally expensive second-order derivative is involved, i.e., $\nabla_{x}^{2}\ell(f_{\theta}(x))$. Fortunately, deep learning and optimization theory [drucker1992improving] allow us to use the first-order derivative to approximate it. We thus choose two well-known first-order gradient regularizations as follows:
@@
where $\lambda_{ir}$ controls the regularization magnitude. IR has been demonstrated improving the generalization and interpretability of DNNs [drucker1992improving, ross2018improving]. Here we show that IR can also improve transferability. Regularizing $\|\nabla_{x}\ell(f(x))\|_{2}$ leads to a smaller curvature $\sigma(\nabla_{x}^{2}\ell(x))$ because the spectral norm of a matrix is upper-bounded by the Frobenius norm, i.e., $\sigma(\nabla_{x}^{2}\ell(x))\leq\|\nabla_{x}^{2}\ell(x)\|_{F}\approx\|\nabla_{x}\ell(f(x))^{T}\nabla_{x}\ell(f(x))\|_{F}\leq\|\nabla_{x}\ell(f(x))\|^{2}$. Here we approximate the second-order derivative using the first-order derivative as suggested in [DBLP:conf/icml/MartensSS12, jakubovitz2018improving]. Consequently, smaller $\sigma(\nabla_{x}^{2}\ell(x))$ indicates better smoothness, thereby improving transferability.
@@
JR has been also correlated with the generalization and robustness [jakubovitz2018improving, yoshida2017spectral, chan2019jacobian, hoffman2019robust, varga2017gradient]. Since JR has a similar form with IR, it is reasonable to presume that they have a similar effect. Formally, we prove that when $\|\nabla_{x}f_{\theta}(x_{i})\|_{F}\rightarrow 0$, we have $\|\nabla_{x}\ell(f_{\theta}(x_{i}))\|\rightarrow 0$ using cross-entropy loss as an example (see Proposition 1). This implies that JR also has a positive effect on model smoothness.
@@ 4.1.2 Regularization in the Weight Space
Recently, researchers paid attention to the gradient regularization on $\theta$, i.e., the parameters (weights) of the network. A recent study [Dherin2022why] proved that the gradient regularizing pressure on the weight space $\|\nabla_{\theta}f_{\theta}(x)\|_{F}$ can transfer to the input space $\|\nabla_{x}f_{\theta}(x)\|_{F}$. Consequently, regularizing gradients w.r.t. the weight space could also implicitly promote the smoothness in the input space. We next explore two representative weight gradient regularization methods as follows:
@@
where $L$ is the original objective function. Similar to IR and JR, we can directly apply the double-propagation technique to implement the regularization term, using the automatic differentiation framework provided by modern ML libraries like PyTorch [paszke2017automatic]. ER has been empirically studied on improving generalization [barrett2021implicit, smith2021on, geiping2022stochastic], while we investigate it from the transferability perspective.
[…]
Recent researches establish SAM as a special kind of gradient normalization [zhao2022penalizing, karakida2022understanding], where $\rho$ represents the regularization magnitude. SAM has been extensively studied recently for its remarkable performance on generalization [foret2021sharpnessaware, andriushchenkoICML22a, chenICLR2022When2022, zhuang2022surrogate, wen2022does, du2022efficient, pmlr-v162-abbas22b]. In this paper, for the first time, we substantiate that SAM also features outstanding superiority in boosting transferability.
@@ 4.2 Gradient Regularization Promotes Smoothness
We train multiple regularized models for each regularization and use coarse-grained parameter intervals [ross2018improving, foret2021sharpnessaware, smith2021on] and restrict the largest regularized magnitude to avoid excessive regularization such that the accuracy of these models are above 90%. This allows for a fair comparison of each regularization method in an acceptable accuracy range. We choose 5 parameters in each interval through random grad and binary choices (See Tab. XI).
@@
Figs. 9 and 10 depict the model smoothness. We can observe gradient regularizations generally result in smoother models (i.e., smaller values) than the baseline where the regularization magnitude is 0 (ST), and large magnitudes correspond to better smoothness. Moreover, input gradient regularizations yield much better smoothness than weight gradient regularizations, particularly on CIFAR-10. This indicates that the transfer effect from weight space to input space suggested in [Dherin2022why] does exist, yet is somewhat marginal and cannot outperform direct regularizations in input space.
@@ 4.3 Trade-off Under Gradient Regularization
One might presume that input gradient regularization (JR, IR) should clearly outperform weight gradient regularization (SAM, ER) in terms of transferability, which agrees with the conclusion that less complexity should exhibit better transferability in Demontis et al. [security19], given that smoothness is positively correlated to the model complexity (as analyzed in Sec. 6, Q3). However, our experimental results counter this:
@@
Therefore, these inconsistent results do not allow us to conclude that better smoothness always leads to better or worse transferability. This finding not only weakens the conclusion in Demontis et al. [security19], but also suggests an overall examination of these training mechanisms for superior surrogates, as they may weigh differently on similarity.
@@ 5.2.1 Incorporating AE generation strategies
Exploiting the generation process of AEs is a popular approach to improve transferability. The most representative methods include momentum iterative (MI) [Dong_2018_CVPR] and diversity iterative (DI) [xie2019improving], and DI is frequently combined with MI and serves as a stronger baseline—diversity iterative with momentum (DIM). Generally, these designs are interpreted as constructing adversarial examples that are less likely to fall into a poor local optimum of the given surrogate. Thus they adjust the update direction with momentum to escape it or apply multiple transformations on inputs to mimic an averagely flat landscape. We conduct experiments to verify that obtaining a better surrogate, which features less and wider local optima (analyzed in Sec. 6, Q3) and general adversarial direction, can further enhance these designs by promoting a larger lower bound.
@@ 5.2.2 Incorporating ensemble strategies
Using multiple surrogates is widely believed to update adversarial examples in a more general direction that benefits transferability. Recently, [gubriLGV2022] proposed constructing the ensemble from a large geometric vicinity (LGV), i.e., first fine-tuning a normally trained model with a high constant learning rate to obtain a set of intermediate models that belongs to a large vicinity, then optimizing AEs iteratively against all these models. We conduct experiments to verify whether SAM and SAM&JR can further enhance this ensemble and analyze how their intrinsic mechanisms correlate with LGV. Specifically, we fine-tune models well-trained with SGD [^3] We refer SGD here to training models without any augmentation or regularization mechanisms, not merely the particular optimizer itself., SAM, and SAM&JR on CIFAR-10 using SGD, SAM, and SAM&JR with a constant learning rate of 0.05.
[…]
@@ 5.2.3 Incorporating other strategies
We take AE generation and ensemble strategies, which are the most well-explored methods, to exemplify the universality of better surrogates. We are aware of other factors that could also contribute to transferability, such as loss object modification [zhao2021on], architecture selection [Wu2020Skip], $l_{2}$-norm consideration [maoP2022TransferAttacks2022], unrestricted generation [hu2022protecting, zhou2023advclip], patch-based AEs [hu2021advhash]. We leave it to our future work to investigate whether they can be mildly integrated with our design.
@@ 6 Analyses and Discussions
Table VI: The Pearson correlation coefficients ($r$) between ASR and accuracy (AccAcc), fooling probability (FPFP), model smoothness (MSMS), and gradient similarity (GSGS). For the coefficient of determination ($R^{2}$) in OLS, $R_{1}^{2}$ depicts the joint effect of $MS$ and $GS$ on ASR and $R_{2}^{2}$ depicts the joint effect of these 4 factors.
@@
Q2: What if the target model is not trained on $\mathcal{D}$? So far, we presume the target models are trained within the normal data distribution $\mathcal{D}$ to lay out our analysis and conclusions, which is also set by default in Demontis et al. [security19]. To verify the generality of the obtained conclusion, we deliver the results here when the data distribution restrictions of target model $\mathcal{G}$ is relaxed. First, we generalize Hypothesis 1 as:
@@
Q3: What is the relationship between model smoothness and model complexity? Model complexity has been investigated in Demontis et al. [security19] to figure out whether it correlates with transferability, which is defined as the variability of the loss landscape, deriving from the number of local optima of the surrogates. Here we correlate the number of local optima with model smoothness. It is obvious that a high-complexity model featuring multiple local optima will cause a larger variance to the loss landscape of AE optimization. Formally, for a feature $x$, the fewer optima in its $\epsilon$-neighbor, the smaller the variance would be. Thus, supposing the loss function used to optimize adversarial points is continuously differentiable, by promoting the smoothness, we can expect fewer optima in the finite $\epsilon$-neighbor space since all the local optima are wider. Consequently, model smoothness negatively correlates with model complexity, i.e., a smoother model indicates lower complexity.
@@
Q4: Does a smaller model generalization gap imply higher adversarial transferability? It is natural to raise this question as the idea of adversarial transferability in a way resembles the concept of model generalization. The former captures the effectiveness of AEs against unseen models, while the latter evaluates the model’s ability to adapt properly to previously unseen data. A concurrent work [wang2022role] attempts to show the positive relationship between model generalization and adversarial transferability, by applying a spectral normalization method constraining the $L_{2}$-operator norm of the surrogate’s weight matrix to empirically demonstrate the increase of both transferability and generalization.
Nevertheless, we emphasize that we cannot over-generalize the alignment between generalization and transferability. It is easy to find some cases that boost generalization yet degrades transferability (e.g., through damaging similarity). For example, the results in Tabs. II, III, XIII and X have shown augmentations that boost generalization could sacrifice transferability. More typically, the strong data augmentations that yield the smallest generalization gaps (see Tab. XII) perform the worst in transfer attacks. In fact, the product of spectral norm of model’s layer-wise weight matrices is a loose upper bound of $\|\nabla_{x}f_{\theta}(x)\|_{2}$ [NEURIPS2020_ab731488, yoshida2017spectral]. When $\|\nabla_{x}f_{\theta}(x)\|_{2}\rightarrow 0$, we have $\|\nabla_{x}f_{\theta}(x)\|_{F}\rightarrow 0$. Utilizing Proposition 1, we also have $\|\nabla_{x}\ell(f_{\theta}(x),y)\|\rightarrow 0$. This suggests that spectral normalization actually benefits transferability by promoting smoothness.
Q5: What is restraining the adversarial transferability of ViT? Recently, ViT [dosovitskiy2021an] receives notable attention for its outstanding performance on various tasks. Meanwhile, some works [shao2022on, Mahmood_2021_ICCV] have demonstrated that ViT models exhibit lower adversarial transferability than CNNs when used as surrogates, and a recent work [naseer2022on] argues that this is because the current attacks are not strong enough to fully exploit the transferability potential of ViT. Complementarily, we provide a different viewpoint that the transferability of ViT may have been restrained by its own training paradigm. ViT is known to have poor trainability [Chen_2021_ICCV, chenICLR2022When2022], and thus generally requires large-scale pre-training and fine-tuning with strong data augmentations. This training paradigm is used by default since its appearance. However, as we discussed in this paper, changing the data distribution could have negative impacts on similarity. Using data augmentation is plausible for accuracy performance, yet not ideal for transfer attacks.